A brick falls off the top of a building. It takes 3 seconds for the brick to hit the ground. How tall is the building?


1
Expert's answer
2021-02-12T10:43:44-0500
"h=\dfrac{1}{2}gt^2=\dfrac{1}{2}\cdot9.8\ \dfrac{m}{s^2}\cdot(3\ s)^2=44.1\ m."